# Generate-on-Graph: Treat LLM as both Agent and KG in Incomplete Knowledge Graph Question Answering

### v1

@@ 7 Conclusion
In this paper, we propose leveraging LLMs for QA under Incomplete Knowledge Graph (IKGQA), which is closer to real-world scenarios, and construct relevant datasets. We propose Generate-on-Graph (GoG), which can effectively integrate the external and inherent knowledge of LLMs. Experiments on two datasets show the superiority of GoG, and demonstrate that an incomplete KG can still help LLMs answer complex questions.
@@
The limitations of our proposed GoG are as follows: (1) We only evaluate GoG on CWQ and WebQSP, which uses Freebase as their background KG. The experimental data may comes from some specific domains. (2) It is possible for LLM to hallucinate in the generation process, which is unavoidable for existing LLMs. (3) There is room for further improvement in performance, as GoG’s performance is lower than that with CoT prompt when KGs are very incomplete.
@@ Appendix A Semantic Parsing Methods Details
{tblr}
width = 0.9colspec = Q[104]Q[837], cell12 = fg=MineShaft, cell21 = fg=MineShaft, vline2 = -, hline1-3 = -, Question & In the nation that spends the Bahamian dollar as currency, what time zone is used?
SPARQL PREFIX ns: <http://rdf.freebase.com/ns/>
SELECT DISTINCT ?x WHERE {
  FILTER (?x != ?c) FILTER (!isLiteral(?x) OR lang(?x) = ” OR langMatches(lang(?x), ’en’))
  ?c ns:location.country.currency_used ns:m.01l6dm .
  ?c ns:location.location.time_zones ?x .
Table 6: An example about "timezone" in the CWQ dataset.
Appendix B Prompt List
@@
{tblr}
width = colspec = Q[90]Q[817]Q[27]Q[27], cell11 = c, cell12 = c, hline1-5 = 1-2, Tasks & Prompt
GoG Instruction Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types:
(1) Search[entity1 | entity2 | …], which searches the exact entities on Freebase and returns their one-hop subgraphs. You should extract the all concrete entities appeared in your last thought without redundant words, and you should always select entities from topic entities in the first search.
(2) Generate[thought], which generate some new triples related to your last thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples.
(3) Finish[answer1 | answer2 | …], which returns the answer and finishes the task. The answers should be complete entity label appeared in the triples. If you don’t know the answer, please output Finish[unknown].
Entities and answers should be separated by tab.
Attention please, entities begin with "m." (e.g., m.01041p3) represent CVT (compound value type) node, and they shouldn’t be selected as the final answers. To find out those entities involved in these event, you could select them as the entities to be searched.
You should generate each step without redundant words.
Here are some examples.
In-Context Few-shot
Question: {Question}
Topic Entity: {List of Topic Entities}
Thought 1:
Filter Relations Please select 3 relations that most relevant to the question and rank them. You should answer these relations in list format directly without redundant words.
Here are some examples.
In-Context Few-shot
Thought: {Thought}
Entity: {Entity}
Relation: {List of Relations}
Answer:
Generate Triples Given the existing triples, please generate some new triples related to your current thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples.
Here are some examples.
In-Context Few-shot
Thought: {Thought}
Known Triples: {Explored Triples}
Generated Triples:
Table 7: Prompts for different tasks used in GoG.
Appendix C Settings for Baselines
@@
Appendix D Statistics of Topic Entities in IKGs
The statistics of topic entities are shown in Table 8, and we drop those samples which have isolated topic entities (topic entity without any neighbor node).
@@ Appendix D Statistics of Topic Entities in IKGs
width = 0.9colspec = Q[108]Q[400]Q[80]Q[80]Q[80]Q[80], column3 = c, column4 = c, column5 = c, column6 = c, cell21 = r=2, cell22 = fg=MineShaft, cell41 = r=2, cell42 = fg=MineShaft, hline1,2,4,6 = -, Dataset & IKG-1 IKG-2 IKG-3 IKG-4
@@
Appendix E Result Analysis
@@
We considered four types of errors: (1) Knowledge Missing. (2) Hallucination. (3) Error before Generate. (4) False Negative. The distribution is shown in Figure 7. It can be found that there are indeed about 26% error are caused by hallucination. Besides, the Figure 6 demonstrates that the performance of the model improves with the number of relevant triples introduced in the Generation step, which means introducing more related triples can alleviate hallucination problems to a certain extent.
Figure 7: The error proportions of GoG under IKG-1 of CWQ.
{tblr}
width = 0.9colspec = Q[185]Q[181]Q[271]Q[271], cells = c, cell13 = fg=MineShaft, cell14 = fg=MineShaft, cell21 = r=2, cell23 = fg=MineShaft, cell24 = fg=MineShaft, cell33 = fg=MineShaft, cell41 = r=2, vline2-4 = 1-5, vline3-4 = 3,5, hline1-2,4,6 = -, Dataset & Models CWQ WebQSP
CKG GPT-3.5 17.7% (55.9) 17.1% (66.6)
GPT-4 4.6% (54.3) 9.7% (65.9)
IKG-1 GPT-3.5 23.7% (57.3) 24.1% (63.4)
GPT-4 8.8% (57.9) 18.4% (70.6)
@@
Table 9: Times of Generate operation in different KG settings. Numbers in brackets represent corresponding Hits@1 score.
Appendix F Case Study
@@
In this section, we present a case analysis to evaluate the utility of GoG, as demonstrated in Table 10. In this case, GoG will first search the neighbors of "Appalachian Mountains" and obtain <Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains>, <Spaulding Mountain, mountain.mountain_range, Appalachian Mountains> and <Old Rag Mountain, mountain.mountain_range, Appalachian Mountains>, then GoG can infer that Appalachian Mountains are also located in North America simply by knowing information about the other three mountains. However, ToG fails on this question once the crucial triple <Appalachian Mountains, location.location.containedby, North America> is missing. In another word, for one-hop questions, GoG can use the neighbors information of the topic entities to predict the tail entities while other methods can not make full use of the neighbors information.
{tblr}
width = colspec = Q[194]Q[748], cell22 = fg=MineShaft, hlines, vline2 = -, Question & Where do the appalachian mountains run?
@@
ToG Reasoning Chains: None
Prediction: I’m sorry, but I need more information to answer your question.
GoG Thought 1: I need to search the Appalachian Mountains to find out where they run.
Act 1: Search[Appalachian Mountains]
Obs 1: Appalachian Mountains, location.contains, Brasstown Bald
Appalachian Mountains, location.contains, Craggy Dome
Appalachian Mountains, location.geolocation, m.02_qbv6
Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains
Spaulding Mountain, mountain.mountain_range, Appalachian Mountains
Old Rag Mountain, mountain.mountain_range, Appalachian Mountains
Thought 2: The Appalachian Mountains contain various locations and are part of geolocations. To find out where they run, I need to generate some new triples based on these given triples and my inherent knowledge.
Act 2:  Generate[where do the Appalachian Mountains run]
Obs 2: Appalachian Mountains, geographical_feature.contains_location, Blue Ridge Mountains
Appalachian Mountains, geographical_feature.start_location, Alabama
Appalachian Mountains, geographical_feature.start_location, North America
Thought 3: The Appalachian Mountains start from North America
Act 3: Finish[Alabama | North America ]
@@
Table 10: Comparison between ToG and GoG.

### v2

@@ 7 Conclusion
In this paper, we propose leveraging LLMs for QA under Incomplete KGs (IKGQA), and construct relevant datasets. We propose Generate-on-Graph (GoG), which can effectively integrate the external and internal knowledge of LLMs. Experiments on two datasets show the superiority of GoG, and demonstrate that an LLMs can be combined with incomplete KGs to answer complex questions.
@@
The limitations of our proposed GoG are as follows: (1) It is possible for LLM to hallucinate in the Generate Action, which is unavoidable for existing LLMs. (2) There is room for further improvement in performance, as GoG’s performance is lower than that with CoT prompt when KGs are very incomplete.
@@
Acknowledgment
This work was supported by Beijing Natural Science Foundation (L243006) and the National Natural Science Foundation of China (No.62376270). This work was supported by the Youth Innovation Promotion Association CAS.
@@ Appendix A Semantic Parsing Methods Details
Appendix B Retrieval Augmented Methods Details
The RA method retrieves relevant paths from the knowledge graph (KG) and uses these paths as context for the large language model (LLM) to generate answers. For instance, ToG employs an LLM to explore the KG, using beam search to select paths related to the question. However, analysis of ToG’s results reveals that approximately 70% of the correct answers come directly from the explored paths, and less than 10% of the correct answers are derived from a combination of the explored path knowledge and the internal knowledge of the LLM. Subsequent experimental results also indicate that under the IKG setting, ToG’s performance is even inferior to that of using the LLM alone. This further demonstrates that such methods do not truly integrate the internal of LLMs and external knowledge of KGs.
Appendix C Prompt List
@@
Appendix D Settings for Baselines
@@
Appendix E Statistics of Topic Entities in IKGs
The statistics of dropped edges are shown in Table 6. Besides, we also ensure that after deleting these crucial triples, the number of neighbor nodes of the topic entities will not be zero. The statistics of topic entities are shown in Table 7, and we drop those samples which have isolated topic entities (topic entity without any neighbor node).
@@ Appendix E Statistics of Topic Entities in IKGs
width = colspec = Q[183]Q[183]Q[183]Q[183]Q[183], columneven = c, column3 = c, column5 = c, hlines, & IKG-20% IKG-40% IKG-60% IKG-80%
CWQ 2.2 4.3 6.4 7.9
WebQSP 6.6 13.9 20.3 27.4
Table 6: The average number of edges deleted for each question under different incompleteness degrees.
{tblr}
width = 0.9colspec = Q[108]Q[400]Q[80]Q[80]Q[80]Q[80], column3 = c, column4 = c, column5 = c, column6 = c, cell21 = r=2, cell22 = fg=MineShaft, cell41 = r=2, cell42 = fg=MineShaft, hline1,2,4,6 = -, Dataset & IKG-20% IKG-40% IKG-60% IKG-80%
@@
Appendix F Compound Value Type (CVT) node
Compound Value Type (CVT) nodes are usually utilized to model events, which could involve start time, end time, location and so on, in KGs. An example of CVT node is illustrated in Figure 5.
Appendix G Result Analysis
G.1 Performance under Generate Action
@@
G.2 Error Analysis
We consider four types of errors: (1) Generate Error, LLMs make error in the Generate Action, such as output wrong entities or "unknown". (2) Decompose Error, LLMs forget the original question after multi-round searching and answer the wrong sub-question in the end. (3) Hallucination, the final answer produced by the LLM is not supported by the evidence in the context (e.g., it lacks some of the constraints), yet the LLM still believes this answer satisfies all the constraints of the question. (4) False Negative, LLMs output the alias name of the ground truth. The distribution is shown in Figure 6. It is evident that the majority of actual errors stem from hallucinations, discounting false negative samples. Moreover, under the IKG setting, there is a higher likelihood of False Negative occurrences due to discrepancies between the answers generated by the Generate Action and the reference answers (for instance, the LLM outputs ’The US’ while the correct answer is "America").
Figure 6: The error proportions of GoG under different datasets and settings.
Models CWQ
CKG IKG-20% IKG-40% IKG-60% IKG-80%
GoG w/GPT-3.5 21.0% (53.8) 33.8% (45.5) 35.9% (52.9) 39.1% (45.2) 39.8% (48.7)
GoG w/Qwen-1.5 24.2% (44.2) 35.5% (42.2) 40.0% (43.5) 46.5% (41.7) 50.4% (43.6)
WebQSP
CKG IKG-20% IKG-40% IKG-60% IKG-80%
GoG w/GPT-3.5 19.3% (63.2) 24.4% (63.5) 26.8% (66.4) 32.5% (57.8) 38.2% (66.4)
GoG w/Qwen-1.5 23.4% (55.9) 28.2% (51.4) 33.9% (57.8) 37.7% (60.2) 49.5% (56.5)
@@ G.2 Error Analysis
Table 8: Ratio of Generate operation in different KG settings. Numbers in brackets represent corresponding Hits@1 score.
Tasks Prompt
GoG Instruction Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity1 | entity2 | …], which searches the exact entities on Freebase and returns their one-hop subgraphs. You should extract the all concrete entities appeared in your last thought without redundant words, and you should always select entities from topic entities in the first search. (2) Generate[thought], which generate some new triples related to your last thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples. (3) Finish[answer1 | answer2 | …], which returns the answer and finishes the task. The answers should be complete entity label appeared in the triples. If you don’t know the answer, please output Finish[unknown]. Entities and answers should be separated by "|". Attention please, entities begin with "m." (e.g., m.01041p3) represent CVT (compound value type) node, and they shouldn’t be selected as the final answers. To find out those entities involved in these event, you could select them as the entities to be searched. You should generate each step without redundant words. Here are some examples. In-Context Few-shot Question: {Question} Topic Entity: {List of Topic Entities} Thought 1:
Filter Relations Please select 3 relations that most relevant to the question and rank them. You should answer these relations in list format directly without redundant words. Here are some examples. In-Context Few-shot Thought: {Thought} Entity: {Entity} Relation: {List of Relations} Answer:
Generate Triples Given the existing triples, please generate some new triples related to your current thought. These new triples may come from your inherent knowledge directly or reasoning from the given triples. Here are some examples. In-Context Few-shot Thought: {Thought} Known Triples: {Explored Triples} Generated Triples:
Verify Triples Given the existing triples please select relevant triples to the question from LLM-generated triples based on your inherent knowledge. Here are some examples. In-Context Few-shotQuestion: {Question} Generated triples: {Generated triples} Answers:
Table 9: Prompts for different tasks used in GoG.
Appendix H Case Study
H.1 Comparison between ToG and GoG under CKG setting
ToG is likely to think compound value types (CVT) are not worthy to further explore and ignore them, as they do not offer information directly. Our GoG can easily solve this problem by expanding subgraph dynamically, that means if there is not enough information provided by the current subgraph, GoG would search one more hop, so the neighbors of CVT nodes is taken into consideration in this way. As illustrated in Table 10, In this case, ToG gets lost and doesn’t retrieve correct information when encounters CVT, "UnName_Entity" represents CVT nodes in the explored paths. On the contrast, our GoG can handle CVT nodes well by further searching.
Question Where did the "Country Nation World Tour" concert artist go to college?
Dropped Crucial Triples None
ToG Reasoning Paths: [(Country Nation World Tour, music.artist.concert_tours, Brad Paisley), (Country Nation World Tour, music.concert_tour.artist, Brad Paisley), (Bachelor’s degree, education.educational_degree.people_with_this_degree, UnName_Entity)], [(UnName_Entity, education.education.institution, University of North Carolina at Chapel Hill), (UnName_Entity, education.education.major_field_of_study, Journalism), (UnName_Entity, education.education.degree, Bachelor’s degree)] Prediction: Based on the given knowledge triplets, the artist who had the "Country Nation World Tour" concert is Brad Paisley. Additionally, it is mentioned that Brad Paisley has a Bachelor’s degree. The knowledge triplets also provide information about the institution he attended, which is the University of North Carolina at Chapel Hill. Therefore, the answer to the question is {University of North Carolina at Chapel Hill}.
GoG Thought 1: I need to find the artist who performed the "Country Nation World Tour" and then find out where they went to college. Action 1: Search[Country Nation World Tour] Observation 1: Country Nation World Tour, concert_tour.artist, Brad Paisley Thought 2: Brad Paisley performed the "Country Nation World Tour", so I need to find out where Brad Paisley went to college. Action 2: Search[Brad Paisley] Observation 2: Brad Paisley, person.education, m.0h3d7qb Brad Paisley, person.education, m.0h3d7qj Brad Paisley, person.education, m.0n1dd_6 Thought 3: m.0h3d7qb, m.0h3d7qj and m.0n1dd_6 could be CVT, I need to further search them. Action 3: Search[m.0h3d7qb | m.0h3d7qj | m.0n1dd_6] Observation 3: Brad Paisley, person.education, m.0h3d7qj m.0h3d7qb, education.institution, John Marshall High School m.0h3d7qj, education.institution, Belmont University m.0h3d7qj, education.major_field_of_study, Music m.0n1dd_6, education.institution, West Liberty University Thought 4: Based on the given observations, Brad Paisley went to Belmont University. This is where the "Country Nation World Tour" concert artist went to college. Action 4: Finish[Belmont University]
Table 10: The comparison between ToG and GoG under CKG setting. The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color.
H.2 Comparison between ToG and GoG under IKG setting
@@
In this section, we present a case analysis to evaluate the utility of GoG, as demonstrated in Table 11. In this case, GoG will first search the neighbors of "Appalachian Mountains" and obtain (Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains), (Spaulding Mountain, mountain.mountain_range, Appalachian Mountains) and (Old Rag Mountain, mountain.mountain_range, Appalachian Mountains), then GoG can infer that Appalachian Mountains are also located in North America simply by knowing information about the other three mountains. However, ToG fails on this question once the crucial triple (Appalachian Mountains, location.location.containedby, North America) is missing. In another word, for one-hop questions, GoG can use the neighbors information of the topic entities to predict the tail entities while other methods can not make full use of the neighbors information.
Question Where do the appalachian mountains run?
@@
ToG Reasoning Paths: None Prediction: I’m sorry, but I need more information to answer your question.
GoG Thought 1: I need to search the Appalachian Mountains to find out where they run. Action 1: Search[Appalachian Mountains] Observation 1: Appalachian Mountains, location.contains, Brasstown Bald Appalachian Mountains, location.contains, Craggy Dome Appalachian Mountains, location.geolocation, m.02_qbv6 Bald Eagle Mountain, mountain.mountain_range, Appalachian Mountains Spaulding Mountain, mountain.mountain_range, Appalachian Mountains Old Rag Mountain, mountain.mountain_range, Appalachian Mountains Thought 2: The Appalachian Mountains contain various locations and are part of geolocations. To find out where they run, I need to generate some new triples based on these given triples and my inherent knowledge. Action 2:  Generate[where do the Appalachian Mountains run] Observation 2: Appalachian Mountains, geographical_feature.contains_location, Blue Ridge Mountains Appalachian Mountains, geographical_feature.start_location, Alabama Appalachian Mountains, geographical_feature.start_location, North America Thought 3: The Appalachian Mountains start from North America Action 3: Finish[Alabama | North America]
@@
Table 11: The comparison between ToG and GoG under IKG setting. The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color.
H.3 Cases of Generate Action
Generate Action is typically invoked when GoG cannot directly obtain an answer from the search results. There are two possible reasons for not being able to directly get an answer from the search results: (1) The correct relation was not selected, or (2) There is a lack of knowledge (the corresponding triples have been deleted). For example, in the case of Table 12, GoG doesn’t select the correct relation "base.biblioness.bibs_location.state" in Search[Montreal], which leads to the answer not appearing in the search results. At this point, GoG uses the internal knowledge of LLMs to generate an answer to the question "which Canadian province is Montreal in" and successfully gets the correct answer "Quebec".
In the case of Table 13, the key triple "Guatemala, location.location.containedby, Central America" is missing, which also leads to the information not being found in the search. At this point, GoG uses the internal knowledge of LLMs along with the retrieved triple "Guatemala, countries.continent, North America" to successfully infer that Guatemala is located in Central America, thereby obtaining the correct answer.
Question Which Canadian Province was the locatiuon for the TV show the Lottery?
Dropped Crucial Triples None
GoG Thought 1: I need to find the location of the TV show ’The Lottery’. Action 1: Search[The Lottery] Observation 1: The Lottery, tv_program.country_of_origin, United States of America The Lottery, tv_program.filming_locations, Montreal The Lottery, tv_program.original_network, m.010pllm8 Thought 2: The Lottery was filmed in Montreal, but I need to identify the Canadian province that Montreal is in. Action 2: Search[Montreal] Observation 2: Hochelaga-Maisonneuve, location.containedby, Montreal Les Bobards, location.containedby, Montreal Mercier, Montreal, location.containedby, Montreal Montreal, travel_destination.tourist_attractions, Canadian Centre for Architecture Montreal, travel_destination.tourist_attractions, Fantasia Film Festival Montreal, travel_destination.tourist_attractions, Marathon Oasis de Montreal Montreal, travel_destination.tourist_attractions, Montreal International Reggae Festival Montreal, travel_destination.tourist_attractions, Montreal Museum of Fine Arts New York Life Insurance Building, location.containedby, Montreal The Pirhana Bar, location.containedby, Montreal Thought 3: Montreal is a city, so I need to find the province it belongs to. Action 3: Generate[which Canadian province is Montreal in] Observation 3: Montreal, location.located_in, Quebec" Quebec, country_part, Canada Thought 4: The answer is the province of Quebec in Canada. Action 4: Finish[Quebec]
Ground Truth Quebec
Table 12: Case for Generate Action in GoG under CKG setting. The correct answers are highlighted with blue color. The Generate Action is highlighted with green color.
Question Which nation has the Alta Verapaz Department and is in Central America?
Dropped Crucial Triples Alta Verapaz Department, location.location.containedby, Guatemala Guatemala, location.location.contains, Alta Verapaz Department Guatemala, location.country.administrative_divisions, Alta Verapaz Department Alta Verapaz Department, location.administrative_division.country, Guatemala Central America, location.location.contains, Guatemala Guatemala, location.location.containedby, Central America
GoG Thought 1: I need to find the nation that contains the Alta Verapaz Department. Action 1: Search[Alta Verapaz Department] Observation 1: Alta Verapaz Department, administrative_area.administrative_area_type, Guatemalan department Alta Verapaz Department, statistical_region.population, g.11bv5vs59b Alta Verapaz Department, statistical_region.population, m.04q66h7 Raxruha, location.containedby, Alta Verapaz Department San Agust0̆0edn Lanqu0̆0edn, location.containedby, Alta Verapaz Department San Pedro Carch0̆0e1, location.containedby, Alta Verapaz Department Santa Catalina la Tinta, location.containedby, Alta Verapaz Department Santa Mar0̆0eda Cahab0̆0f3n, location.containedby, Alta Verapaz Department Thought 2: I found that Alta Verapaz Department is a Guatemalan department, now I need to confirm if Guatemala is in Central America. Action 2: Search[Guatemala] Observation 2: Guatemala, countries.continent, North America Guatemala, country.languages_spoken, Spanish Language Thought 3: The information says Guatemala is in North America, not Central America. This may be due to the definition of C̈entral American̈ot being strictly a continent but a region. I need to verify if Guatemala is considered a part of Central America based on the given topic entity. Action 3: Generate[Central America includes Guatemala] Observation 3: Alta Verapaz Department, country.administrative_area_of, Guatemala Central America, region.includes, Guatemala Guatemala, country.geographic_region, Central America North America, continent.includes, Central America Thought 4: Based on the new triples, I can confirm that Guatemala, which contains the Alta Verapaz Department, is indeed in Central America. Action 4: Finish[Guatemala]
Ground Truth Guatemala
Table 13: Case for Generate Action in GoG under IKG setting. The correct answers are highlighted with blue color. The Generate Action is highlighted with green color.
H.4 Bad Cases of GoG
Hallucination
The case in Table 14 demonstrates an example of GoG, where GoG mistakenly inferred that the date "2012-01-01" was when the team won the championship, while in fact, this date marks the beginning of Larr Baer’s leadership of the team.
Question When was the last time a team led by Larr Baer win the World Series?
Dropped Crucial Triples None
GoG Thought 1: I need to find the team led by Larry Baer and then find out when they last won the World Series. college. Action 1: Search[Larry Baer] Observation 1: Larry Baer, board_member.leader_of, m.011zsc4_ Thought 2: I need to find out the team led by Larry Baer and then find out when they last won the World Series. Action 2: Search[m.011zsc4_] Observation 2: Larry Baer, board_member.leader_of, m.011zsc4_ San Francisco Giants, organization.leadership, m.011zsc4_ m.011zsc4_, leadership.from, 2012-01-01 Thought 3: The San Francisco Giants is the team led by Larry Baer and they last won the World Series in 2012. Action 3: Finish[2012-01-01]
Ground Truth 2014 World Series
Table 14: Case for hallucination in GoG. The wrong answers are highlighted with red color.
Generation Error
The case in Table 15 demonstrates an example of GoG, where GoG make an error in generating the mascot for Syracuse University Athletics. It mistakenly identified "Orangeman" as the mascot of the team, but in reality, the team’s mascot is Otto the Orange. "Orangeman" is actually the name used to refer to the team.
Question What is Syracuse University Athletic’s mascot?
Dropped Crucial Triples None
GoG Thought 1: I need to find the mascot of Syracuse University Athletics." Action 1: Search[Syracuse University Athletics] Observation 1: Syracuse University Athletics, organization.headquarters, m.0_gdsbw Syracuse University Athletics, organization.parent, m.0105kgwd m.0_gdrcn, award_honor.award_winner, Syracuse University Athletics m.0_y3csk, award_honor.award_winner, Syracuse University Athletics Thought 2: There is no information about the mascot of Syracuse University Athletics, so I need to generate more triples based on these given triples and my inherent knowledge. Action 2: Generate[What is the mascot of Syracuse University Athletics] Observation 2: Syracuse University Athletics, athletics_brand.mascot, Orangemen Thought 3: The mascot of Syracuse University Athletics is Orangemen. Action 3: Finish[Orangemen]
Ground Truth Otto the Orange
Table 15: Case for generation error in GoG. The wrong answers are highlighted with red color.
Decompose Error
The case in Table 16 demonstrates an example of GoG. In this example, during the process of decomposing the problem, GoG forgets that the initial goal is to find the team coached by Pablo Laso. Instead, in the final thought, the objective shifts to finding the country where this team is located. This type of situation is likely to become more frequent as the number of search iterations increases and the context length grows longer.
Question What basketball team in Spain is Pablo Laso head coach for?
Dropped Crucial Triples None
GoG Thought 1: I need to find the basketball team that Pablo Laso is the head coach for in Spain. Action 1: Search[Pablo Laso] Observation 1: Pablo Laso, basketball_coach.team, Real Madrid Baloncesto Pablo Laso, pro_athlete.teams, m.012n525v Real Madrid Baloncesto, basketball_team.head_coach, Pablo Laso Thought 2:  Pablo Laso is the head coach for Real Madrid Baloncesto, so I need to find out which country Real Madrid Baloncesto is located in. Action 2: Search[Real Madrid Baloncesto] Observation 2: Real Madrid Baloncesto, sports_team.arena_stadium, Palacio de Deportes de la Comunidad de Madrid Real Madrid Baloncesto, sports_team.arena_stadium, Raimundo Saporta Pavilion Real Madrid Baloncesto, sports_team.location, Madrid Real Madrid Baloncesto, sports_team.location, Spain Real Madrid Baloncesto, sports_team.venue, m.0wz2_rd Real Madrid Baloncesto, sports_team.venue, m.0wz2wth Thought 3:  Real Madrid Baloncesto is located in Madrid, Spain. Action 3: Finish[Spain]
Ground Truth Real Madrid Baloncesto
Table 16: Case for decompose error in GoG. The wrong answers are highlighted with red color.